OpenStudy (anonymous):

Simplify completely quantity 5 x squared plus 9 x minus 2 over quantity x squared plus 12 x plus 20 times quantity x squared plus 17 x plus 70 over quantity 15 x minus 3

OpenStudy (johnweldon1993):

\[\large \frac{5x^2 + 9x - 2}{x^2 + 12x + 20} \times \frac{x^2 + 17x + 70}{15x - 3}\] Like that right?

OpenStudy (anonymous):

yes that looks right

OpenStudy (johnweldon1993):

Well this can be simplified a lot through factorization ...as in So using factorization...how can we simplify (lets do...) x^2 + 17x + 70 ?

OpenStudy (anonymous):

(x+7)(x+10)?

OpenStudy (johnweldon1993):

Perfect...alright so we'll put that to the side for now...how about x^2 + 12x + 20 ?

OpenStudy (anonymous):

(x+2)(x+10)

OpenStudy (johnweldon1993):

Good again...now 15x - 3 ? what number can be factored out?

OpenStudy (anonymous):

3(5x-1)

OpenStudy (johnweldon1993):

Doing awesome! and finally 5x^2 + 9x - 2 ?

OpenStudy (anonymous):

gosh I hate these higher coefficients give me a min

OpenStudy (johnweldon1993):

Of course :)

OpenStudy (anonymous):

(5x-1)(x+2) sorry finally haha

OpenStudy (johnweldon1993):

Perfect! and I was gonna say you do know the fast way to do those right? But alright good :)

OpenStudy (johnweldon1993):

So now...altogether we have \[\large \frac{(5x - 1)(x + 2)}{(x + 10)(x + 2)} \times \frac{(x + 10)(x + 7)}{3(5x - 1)}\]

OpenStudy (johnweldon1993):

Since this is multiplication...we can just combine these over 1 fraction... \[\large \frac{(5x - 1)(x + 2)(x + 10)(x + 7)}{(x + 10)(x + 2)3(5x - 1)}\]

OpenStudy (johnweldon1993):

Now we can begin to cancel...what is on the top as well as the bottom...I see a bunch \[\large \frac{\cancel{(5x - 1)}(x + 2)(x + 10)(x + 7)}{(x + 10)(x + 2)3\cancel{(5x - 1)}}\]

OpenStudy (johnweldon1993):

\[\large \frac{\cancel{(5x - 1)}\cancel{(x + 2)}(x + 10)(x + 7)}{(x + 10)\cancel{(x + 2)}3\cancel{(5x - 1)}}\]

OpenStudy (johnweldon1993):

\[\large \frac{\cancel{(5x - 1)}\cancel{(x + 2)}\cancel{(x + 10)}(x + 7)}{\cancel{(x + 10)}\cancel{(x + 2)}3\cancel{(5x - 1)}}\]

OpenStudy (johnweldon1993):

So all that is left is \[\large \frac{x + 7}{3}\]

OpenStudy (anonymous):

Sweet! I actually understand that. The quiz I needed it for is long finished but I'm really glad I got to go back through it. Thank you so much! You've been a great help. :)

OpenStudy (johnweldon1993):

Anytime hun :)